Log(x+2) - log(x-1) = logx - log(x+3)
Answer by ewatrrr(24785)

 
Hi,
Log(x+2) - log(x-1) - logx + log(x+3)= 0


Log (x+2)(x+3)/x(x-1) = 0
(x+2)(x+3)/x(x-1) = 10^0 = 1
(x+2)(x+3)= x(x-1)
x^2 + 5x + 6 = x^2 - x
6x= -6
x = -1 Extraneous solution
NO solution